### v2

# Chirality-induced Spin-Orbit Coupling and Spin Selectivity

Abstract
We show that a spinor traveling along a helical path develops a spin-orbit coupling because of the curvature of the path. We then estimate the magnitude of this effective geometric spin-orbit interaction for structures that showcase chirality-induced spin selectivity (CISS). We find that this chirality-induced spin-orbit coupling ($\chi$-SOC), in conjunction with broken time-reversal symmetry, may be an important ingredient for the microscopic underpinning of the CISS phenomenon.
spin-orbit interaction, curvilinear coordinates, spin selectivity
Introduction – Spin-orbit coupling (SOC) is a fundamental relativistic phenomenon arising from the coupling between the spin and orbital degrees of freedom of a spinful particle [1]. In atomic systems, it is typically smaller the lighter the atoms. The spin itself can be manipulated by coupling it to a magnetic field. Therefore, it was a complete surprise that chiral molecules made of light atoms and hence weak SOC, display a spin response (selectivity) in the absence of any external magnetic field [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This phenomenon has been called chirality-induced spin selectivity (CISS) and it has triggered a large amount of research in physics, chemistry, and biology, also in view of the broad spectrum of potential applications it may offer [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
There is agreement that SOC must play a key role in determining the CISS effect  [31, 32, 33, 34, 35, 36, 37, 38, 39] as well as time-reversal symmetry breaking via an applied voltage in transport junctions or by decoherence. [36, 40, 41, 42, 43, 44, 45] However, while time-reversal symmetry breaking is relatively easy to account for, in view of the way experiments are performed, the origin of a possible non-negligible SOC is still under debate.  [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 39] and the coupling to additional degrees of freedom has also been discussed. [49, 50, 56, 51, 57]
Furthermore, on the side of first-principles calculations, there is no full agreement concerning the orders of magnitude of the spin polarization, [58, 59, 60, 42, 61, 62] so that the ultimate origin of the CISS effect remains under debate. In particular, a recent study [62] based on a fully relativistic density functional theory methodology combined with the Landauer-Büttiker approach has suggested the need to include geometric terms in the SOC to achieve closer agreement with experimental trends.
Figure 1: Schematic representation of the model system. A helical tube is considered, which is subsequently mapped onto a one dimensional helical path. The full spinor wave function consists of a spatial component $\Phi(\phi)$ and a general spinor on which an SU(2) rotation operator is acting. The angle $\phi$ parametrizes the helix arc length, $\kappa=\pm 1$ labels the helicity, and $\vec{n}$ is the spin rotation axis. The eigenvalues of the Hamiltonian in Eq. (5) are schematically shown on the right panel as a function of the angular momentum $l$ for the two spin branches (“+” and “--”), cf. Eq. (9).
It is, therefore, very appealing to see spin-orbit coupling emerging from a general geometric principle. In Refs.  [63, 64, 65] such a geometric SOC was derived. Shitade and Minamitani [63] started from the Dirac Lagrangian density in a curved space-time to arrive at an SOC expression proportional to the curvature of a helical system. This SOC includes the product of the projection of the Pauli spin vector $\vec{\sigma}$ in the direction of the helix binormal vector $\vec{B}$ (using a Frenet-Serret basis), and the linear momentum $p_{s}$ of the electron along the helix: $(\vec{\sigma}\cdot\vec{B})p_{s}$. An estimate of the coupling strength was obtained to be approximately 160 meV, which is far stronger than any atomic SOC of light atoms. In addition, there seems to be a potential issue with the approach of Shitade and Minamitani, which the authors also acknowledge: the final results can be different dependent on whether the thin-layer quantization is performed before or after the Foldy-Wouthuysen transformation to obtain the non-relativistic limit of the Dirac equation. In fact, Yu [64, 65] exploited the relativistic equivalence of a curved space-time manifold and a noninertial system to obtain a different result, in terms of the local normal vector $\vec{N}$: $(\vec{N}\times\vec{p})\cdot\vec{\sigma}$. An estimate of the coupling constant yielded in this case 0.2 meV, but for a reference polymeric system with a much larger radius and pitch than, e.g., DNA. We also remark that the above issue only appears when the Dirac equation is taken as a starting point: using directly the Pauli equation [66, 67] or the spin-independent Schrödinger equation [68, 69] does not lead to any inconsistencies when using the thin-layer approximation.
Unlike previous studies, here we show, using a spin-independent Hamiltonian, that the dynamics of a spinful particle along a helical path naturally develops a purely kinetic effective SOC, even if the particle does not experience any other potential (besides a spin-independent confinement potential transverse to a helical path). We find that this chirality-induced SOC, which we denote as $\chi$-SOC, is substantial for systems that currently show the CISS effect. We suggest that, together with the breaking of time-reversal symmetry (originating, e.g., from the external bias applied in the experiments), this $\chi$-SOC provides a simple way to generate an SOC in chiral systems.
Model Hamiltonian – We consider a simple Hamiltonian of an electron on a curved path, in particular, on an infinite helical tube with finite cross section (see Fig. 1). Only the kinetic energy of the particle is considered, together with a potential leading to confinement in the transverse direction to the helical curve. The choice of the confinement may play a role in determining the particular form of the effective 1d Hamiltonian to be obtained in the next step, but we are limiting ourselves to the simple choice of a confinement potential with SO(2) symmetry, e.g., of the form $V_{\lambda}(q_{1},q_{2})=(1/2)\lambda^{2}(q_{1}^{2}+q_{2}^{2})$, where $q_{1},q_{2}$ are local transverse coordinates along the helical curve. However, no initial spin-orbit interactions are taken into account. As originally shown by da Costa [68], and more recently by Geyer et al. [38, 70] within a rigorous space-adiabatic framework, one can decouple longitudinal and transverse degrees of freedom to map the 3d structure of the helical tube on a one-dimensional (infinite) helical path --a similar approach has also been used, e.g., in Refs. [71, 66]. Since we only consider the kinetic energy and a scalar confinement, the Hamilton operator adopts a very simple form (see the Supplementary Information for its derivation):
$\hat{H}_{0}=-\frac{\hbar^{2}}{2mL^{2}}\{\frac{\partial^{2}}{\partial\phi^{2}}+\frac{\rho R}{4}\}.$ (1)
Here, $\phi$ is an angular variable related to the arc length, $s$, of the helix by the relation $\phi=s/L$, $L=\sqrt{R^{2}+(b/2\pi)^{2}}$ being the length of a single turn of a helix with radius $R$ and pitch $b$. The helix curvature is defined as $\rho=R/L^{2}$. The second term in Eq. (1) is a quantum geometric potential already obtained by da Costa, [68] but also in other studies. [71, 66, 69]
Consider now the general representation of a spinor wave function on the helix whose Hamiltonian is described by Eq. (1):
$\vec{\Psi}(\phi)=\exp{-i\kappa\frac{\phi}{2}\vec{n}\cdot\vec{\sigma}}\vec{\chi}\otimes\Phi(\phi)=\mathcal{U}(\phi)\vec{\chi}\otimes\Phi(\phi).$ ) . (2)
Due to the absence of SOC in Eq. (1) the spin and spatial components are separable. The 2-component spinor $\vec{\chi}$ does not need to be specified at this stage, its components will be calculated later on. Notice that the spin rotation is tied to the space frame of the helix. The unitary operator acting on the spinor $\vec{\chi}$ induces a spin rotation around $\vec{n}$ while the electron moves along the helix (this is similar to the action of a quantum gate on a qubit, with the helix playing the role of the “quantum gate”). The parameter $\kappa=\pm 1$ accounts for a change from a right-handed to a left-handed helix, since the sign of $\phi$ changes in this case.
The spatial part $\Phi(\phi)$ can be written as a linear combination of “plane wave” solutions with (real valued) angular momentum $l$ as:
$\Phi(\phi)=\int_{-\infty}^{\infty}\frac{dl}{\sqrt{2\pi}}\,A_{l}e^{il\phi},$ (3)
with $\int_{-\infty}^{\infty}\frac{dl}{{2\pi}}\,|A_{l}|^{2}=1$. We can also restrict the integration to positive values of $l$ by introducing the index $s=\mathrm{sgn}(l)=\pm 1$: $e^{il\phi}\rightarrow e^{is|l|\phi}$. Acting with the Hamiltonain Eq. (1) on the wave function Eq. (2), and defining $E_{0}=\hbar^{2}/2mL^{2}$ as a characteristic energy scale of the problem, we obtain:
$\hat{H}_{0}\vec{\Psi}(\phi)$ $=$ $E_{0}\mathcal{U}(\phi)\{(-i\frac{\partial}{\partial\phi})^{2}-\kappa(\vec{n}\cdot\vec{\sigma})(-i\frac{\partial}{\partial\phi})$
$+$ $\frac{1}{4}(\vec{n}\cdot\vec{\sigma})^{2}-\frac{\rho R}{4}\}\vec{\chi}\otimes\Phi(\phi)$
$=$ $E_{0}\mathcal{U}(\phi)\{p_{\phi}^{2}-\kappa(\vec{n}\cdot\vec{\sigma})p_{\phi}$
$+$ $(\frac{1}{4}-\frac{\rho R}{4})\}\vec{\chi}\otimes\Phi(\phi).$
In the second equality, we have introduced the angular momentum operator $p_{\phi}=-i\partial/\partial\phi$ and used the result $(\vec{n}\cdot\vec{\sigma})^{2}=1$. Therefore, we can introduce a new effective Hamiltonian as:
$\hat{H}=E_{0}\{[p_{\phi}^{2}+\frac{1}{4}(1-\rho R)]\mathds{1}_{2\times 2}-\kappa(\vec{n}\cdot\vec{\sigma})p_{\phi}\},$ (5)
where we stress the fact that the kinetic energy operator and the correction leading to a geometric potential are both diagonal in spin space. These results show that it is possible to derive an effective spin-orbit coupling for an electron moving on a curvilinear path, even if no previous SOC was present. The key result is that the geometric phase accumulated by the spin during its motion leads to an effective interaction between the spin and the orbital degrees of freedom, which can be interpreted as a chirality-induced spin-orbit coupling term:
${(\vec{L}}\cdot{\vec{S})}_{\mathrm{\chi-SOC}}\equiv(2E_{0}/\hbar)\kappa(\vec{n}\cdot\vec{S})p_{\phi}.$ (6)
Notice that the obtained $\chi$-SOC has a purely kinetic origin and its strength is controlled by the energy scale $E_{0}$. For a DNA helix with $R=1\,nm$ and $b=3.4\,nm$, one estimates $E_{0}\approx 30\,meV$, which is larger by a factor 3 to 4 than the atomic SOC of light elements.  [^1] However, note that in this estimate we have used the free electron mass $m$. In a more accurate calculation, an effective mass $m^{*}$ should be employed. The obtained geometric SOC is clearly time-reversal invariant, and it can also be rewritten as an SU(2) “pseudo-gauge field” by completing squares in Eq. (5):
$\hat{H}$ $=$ $E_{0}(p_{\phi}\mathds{1}_{2\times 2}-e\mathcal{A})^{2}-\frac{\hbar^{2}}{8m}\rho^{2},$ (7)
$\mathcal{A}$ $=$ $\frac{\kappa}{2e}\vec{n}\cdot\vec{\sigma}.$ (8)
This approach leverages the properties of SU(2) rotations to capture the evolution of the spin state in a curved trajectory. Note that this contribution would also be present on a circle ($b=0$), although in this case the angular momentum variable would be quantized due to the periodicity condition $\Phi(\phi+2\pi)=\Phi(\phi)$, but it would trivially vanish as $R\rightarrow\infty$, i.e., in the limit of a straight line.
The eigenvalues of Eq. (5) can be easily found:
$\tilde{E}^{\kappa}_{\pm,l}=l^{2}+\frac{1}{4}(1-\rho R)\pm\kappa l=(l\pm\frac{\kappa}{2})^{2}-\frac{\rho R}{4},$ (9)
where $\tilde{E}^{\kappa}_{\pm,l}=(E^{\kappa}_{\pm,l})/E_{0}$. This represents two parabolas shifted horizontally from each other by $\kappa$ (see schematic in Fig. 1). The eigenvalues as a function of $l$ yield two spin branches (“+” and “--”) for positive $l$ (or $s$ = 1), and another two spin branches for negative $l$. Due to time-reversal symmetry, the relation $\tilde{E}^{\kappa}_{+,l}=\tilde{E}^{\kappa}_{-,-l}$, so that Kramer’s theorem holds, as expected.
The corresponding spinor eigenfunctions $\vec{\chi}$ can be obtained in terms of the components of the vector $\vec{n}$, which we parametrize in general using two angles $\alpha,\beta$ as $\vec{n}=(\sin\alpha\cos{\beta},\sin\alpha\sin{\beta},\cos\alpha)$ ). In this way, we get the following:
$\vec{\chi}_{+}=e^{i\beta/2}\left(\begin{array}[]{l}\sin{(\alpha/2)}e^{-i\beta/2}\\ -\cos{(\alpha/2)}e^{i\beta/2}\end{array}\right),$ (12)
and
$\vec{\chi}_{-}=e^{-i\beta/2}\left(\begin{array}[]{l}\cos{(\alpha/2)}e^{-i\beta/2}\\ \sin{(\alpha/2)}e^{i\beta/2}\end{array}\right).$ ) . (15)
Another pair of eigenvectors is obtained for $s=-1$ simply by replacing $e^{is|l|\phi}\rightarrow e^{-is|l|\phi}$.
[…]
$\hat{\nu}_{\phi}=\frac{E_{0}L}{\hbar}\{-2i\frac{d}{d\phi}\mathds{1}_{2\times 2}-(s\kappa)\,\vec{n}\cdot\vec{\sigma}\}.$ (16)
Using the latter expression, the charge current is obtained as:
$j^{\pm,\kappa,s}_{c}$ $=$ $\frac{e\hbar}{mL^{2}}s\int_{0}^{\infty}\frac{dl}{{2\pi}}\,|A_{l}|^{2}(l\pm\frac{\kappa}{2})$ (17)
which gives a total charge current of $j^{\kappa,s}_{c}=(2e\hbar/mL^{2})s\int_{0}^{\infty}\frac{dl}{{2\pi}}\,|A_{l}|^{2}l$ [^2] Notice that for systems with a very small pitch, transport pathways along the helical axis may interfere (due to tunneling) with those following a helical path, and eventually lead to a reduction of chirality-induced effects. This is, indeed the case of helicene, as shown in Ref. [79] using first-principles calculations.
The difference of these currents for a given propagation direction, e.g., $s=1$, yields $(e\hbar/mL^{2})s\kappa$, which is proportional to the helicity $\kappa$, and thus changes sign upon a mirror inversion operation. The fact that this difference does not vanish indicates that the spins in the (+) and (--) states propagate with different velocities. This, in particular, allows us to define a spin polarization ($\mathrm{SP}$) of the charge current:
$\mathrm{SP}=(j^{+,\kappa,s}_{c}-j^{-,\kappa,s}_{c})/j^{\kappa,s}_{c}=\kappa/(2\int_{0}^{\infty}\frac{dl}{{2\pi}}\,|A_{l}|^{2}l).$ (18)
We remark that in a real system, finite-size quantization will lead to discrete values of the $l$ quantum number and the integrals will become summations: $\int_{0}^{\infty}\frac{dl}{2\pi}\rightarrow\sum_{l=1}^{\infty}$.
In the special case of a single $l$-mode contributing to the summation, we can make a rough estimate of the spin polarization by assuming $A_{l}\sim\delta(l-l_{0})$, so that $\mathrm{SP}=\kappa/(2l_{0})$, which for $l_{0}=1$ gives a 50% polarization for $\kappa=1$. This is of the same order of magnitude of measured spin polarizations in, e.g., DNA [2]. Notice, however, that this result already assumes broken time-reversal symmetry, since there would be another mode with $l=-l_{0}$ yielding the opposite polarization. Another example, using a Gaussian profile is discussed in the Supplementary Information.
More realistic estimates would require the formulation of a full spin transport problem with inclusion of scattering effects at, e.g., substrate-molecule interfaces, and those due to interactions with, e.g., vibrational degrees of freedom. [56, 41] We also point out that even though the magnitude of spin polarization may be different, the phenomenon we predict is present irrespective of whether the transport mechanism is ballistic or hopping [73]. The reason is that even in the latter case the helix would create an effective spin-orbit coupling of the type we derive.
In a similar way, the spin currents can be calculated, yielding:
$j^{\pm,\kappa,s}_{spin}(\alpha)$ $=$ $\mp\frac{\hbar^{2}}{2mL}s\int_{0}^{\infty}\frac{dl}{{2\pi}}\,|A_{l}|^{2}(l\pm\frac{\kappa}{2})\cos{\alpha},$ (19)
which leads to the total spin current $j^{\kappa,s}_{spin}(\alpha)=-(\hbar^{2}/2mL)\kappa s\cos{\alpha}$ ). Using the above helix parameters of DNA, we can obtain an estimate of the coefficient of the spin current $\hbar^{2}/2mL=0.332\,eV\,nm$.
Notice that the symmetries of the spin current are: $j^{\pm,-\kappa,s}_{spin}(\alpha)=-j^{\pm,\kappa,s}_{spin}(\alpha)$, and $j^{\pm,-\kappa,-s}_{spin}(\alpha)=j^{\pm,\kappa,s}_{spin}(\alpha)$, i.e., changing the chirality changes the sign of the spin current, while a change in chirality together with time-reversal ($s\to-s$) leaves the spin current invariant. A non-zero spin current only emerges if time-reversal symmetry is broken, otherwise any contribution for $+s$ will get canceled by a similar term with $-s$.
In a geometric picture, the chirality parameter $\kappa$ should be related to the helix torsion in a suitable dimensionless quantity. It is also interesting to note that our results have a qualitative resemblance to the analytical model presented in Ref. [75], which, however, introduces a Rashba spin-orbit interaction in the Hamiltonian from the very beginning. Moreover, the angle $\alpha$ in our case parameterizes the spin rotation vector, while in Ref. [75] it is related to the strength of the Rashba spin-orbit coupling.
Conclusions– In conclusion, we have shown that a spinful particle traveling along a helical path naturally develops an effective SOC, even without an intrinsic SOC. This chirality-induced SOC ($\chi$-SOC) is much stronger than the typical relativistic SOC of light atoms, thus providing an additional source of spin polarization. Our results suggest a possible strong additional contribution to the chirality-induced spin selectivity observed experimentally in chiral organic and inorganic materials. In future work, it would be interesting to address issues like the temperature and length dependence of this effect, which would require a spin transport calculation, including the interaction with dynamical degrees of freedom such as linear or chiral phonons. Regardless of the effect, our work shows that geometric effects introduced by a chiral structure lead to a novel type of SOC, otherwise absent in traditional atomic systems. It is worth mentioning that our model Hamiltonian can be formally adapted to different spin transport setups, including two-terminal measurements, as in break junctions, [76] as well as more complex setups addressing a ”transverse” CISS effect. [77, 78] We leave this for future studies.
Acknowledgements – The authors thank Matthias Geyer for very fruitful discussions. M. D. acknowledges funding by the Alexander von Humboldt Stiftung through the 2024 Humboldt Research Award. R.G. and G.C. acknowledge the support of the German Research Foundation (DFG) within the project Theoretical Studies on Chirality-Induced Spin Selectivity (CU 44/55-1), and by the transCampus Research Award Disentangling the Design Principles of Chiral-Induced Spin Selectivity (CISS) at the Molecule-Electrode Interface for Practical Spintronic Applications (Grant No. tCRA 2020-01) and Programme trans-Campus Interplay between vibrations and spin polarization in the CISS effect of helical molecules (Grant No. tC2023-03).
Author contribution – M.D. and R.G. contributed equally to this work.